A recent college graduate buys a new car by borrowing $18,000 at 7.2%, compounded monthly, for 5 years. She decides to pay $3Please round to 2 decimal places.
Answer #1

As per the answers of the first part, EMI has already been calculated, which is appx $358.12

(b). If her EMI of 60 months is $ 358.12, that means she has to pay a total of

358.12 * 60 = $21487.2 in five years.

Now, as per the question, if she pays $369 per month, then she will pay for 21487.2/369 = 58.23 months

Appx for 58.2 months

(c). If she pays 369 per month over the entire loan period, then she has to pay 369*60 = $22140

(d) If she only pays her EMI, then her total amount would have 358.12*60 = $21487.2

(e) If she pays $369 per month instead of her EMI of $358.12, then approximately, her 1.8 EMIs (as solved in the part b above) will be saved

that means, 358.12 * 1.8 = $644.16

approximately, she will save a total of $644.16 y paying $369 per month.